Space Stations

NASA ISS On-Orbit Status 23 February 2015

By Marc Boucher
Status Report
February 24, 2015
Filed under , ,
NASA ISS On-Orbit Status 23 February 2015
NASA astronaut Terry, Virts Flight Engineer of Expedition 42 is seen working to complete a cable routing task while near the forward facing port of the Harmony module on the International Space Station. Virts and fellow astronaut Barry "Butch" Wilmore completed a 6-hour, 41-minute spacewalk routing more than 300 feet of cable as part of a reconfiguration of the station to enable U.S. commercial crew vehicles under development to dock to the space station in the coming years. Credit: NASA.

Today: Drain Brain: Cristoforetti set up the Human Research Facility (HRF) Pulmonary Function System (PFS) and initiated hardware warm-up in support of Drain Brain Plethysmography measurement.
The crew verified the Ultrasound 2 and Video Power Converter configuration, set up a video camera and performed an ultrasound scan with guidance from the Drain Brain ground team. On Earth, blood flows down from a person’s brain back toward the heart thanks in part to gravity, but very little is known about how this flow happens without gravity’s effects. Many crewmembers report headaches and other neurological symptoms in space, which may be related to the absence of gravity acting on blood flowing through the veins. Drain Brain uses a special neck collar to measure blood flow from the brain, to help researchers understand which physical processes in the body can compensate for the lack of gravity to ensure blood flows properly.

Ansio Tubule: Cristoforetti retrieved an Aniso Tubule Sample from Minus Eighty degree Laboratory Freezer for ISS (MELFI) and attached samples on the Saibo rack front to expose to cabin light. After 6 – 24 hours the crew will insert the sample into the Cell Biology Experiment Facility (CBEF). Experimenters will examine growth modifications of Arabidopsis hypocotyls and the changing orientation of cortical microtubule under microgravity conditions in space. Scientists will further observe changes in dynamics of cortical microtubules and microtubule-associated proteins by observing Arabidopsis hypocotyls and microtubule-associated proteins with a fluorescence microscope.

Minus Eighty Degree Laboratory Freezer for ISS (MELFI): Virts powered off MELFI and removed the cold box front panel. The crew next inspected the Brayton Machine to determine if any condensation is present and evaluated electrical connectors for corrosion and downlinked photos to the ground teams. Braycote 601 was applied to electrical connectors for corrosion protection, the front panel reinstalled MELFI powered back on.

Urine Processing Assembly (UPA) Status: As a result of last Tuesday’s high pressure sensor (P11) fault, today the crew performed troubleshooting and recovery steps on the UPA by first removing the top cover from the Fluids Control Pump Assembly (FCPA) and performing an inspection. During the inspection, brine was found on the FCPA surface, which led to the decision to replace the unit. UPA has since been repowered. Later this evening, Flight Controllers plan to perform a process cycle.

Extravehicular Activity (EVA) Preparation: Today, Wilmore removed Extravehicular Mobility Unit (EMU) 3005 from the Forward EMU Don/Doff Assembly (EDDA) and installed EMU 3010 in order to perform a fan dry out on the suit. Once the dry out activity was completed, he removed EMU 3010 from the EDDA and installed 3005 back into the EDDA. Later, all three USOS crew members conducted a review of procedures associated with EVA #30 to complete International Docking Adapter (IDA)-1 cable install and perform Space Station Remote Manipulator System (SSRMS) Lee A Lube. EVA #30 is scheduled for Wednesday, February 25th with Joint Airlock Egress occurring at ~6:15am CST.

Mobile Servicing System (MSS) Operations: Today, the Robotics Ground Controllers powered up the MSS and stowed the Special Purpose Dexterous Manipulator (SPDM) on Mobile Base System (MBS) Power Data Grapple Fixture (PDGF) 2. They then maneuvered the Space Station Remote Manipulator System (SSRMS) into position for the start of USOS EVA #30.

At the end of rigidizing its Latching End Effector (LEE) on MBS PDGF2, the SPDM safed itself. The Robotics Ground Controllers subsequently performed the necessary recovery steps and latched and mated the SPDM LEE to the PDGF. This safing is thought to have been caused by the SSRMS and SPDM being in a hard to backdrive configuration for the PDGF capture and will be reported as a recurrence to a previous anomaly. Both the SSRMS and the SPDM are in a nominal configuration for the lubrication of SSRMS LEE A during Wednesday’s EVA.

Today’s Planned Activities
All activities have been completed unless otherwise noted.

HRF – saliva sampling
HRF – place samples into MELFI
Testing [??? ??]
DB- equipment setup and heating activation
Info d/l from Glisser-M video complex (term)
EMU dryout
ARED quarterly inspection
Activation of HRF1 Ultrasound 2
DB – experiment run, part 1
CASCADE. Manual mixing in bio reactor
DB-USND-operations for the experiment
Inspect cnctr J4 FCPA of UPA system
Provide access to P11 sensor in FCPA of UPA
BIOPLENKA. Kriogem-03 tear down
HRF- equipment stow
ANISO – retrieval of samples from MELFI
ANISO – install packages with samples on CBEF door
Psychological assessment program run (WinSCAT)
Disconnect failed P11 sensor and install in its place a by-pass in FCPA of UPA
EMU cartridge regeneration (term)
WHS – replace urine receptacle and filter insert
[???? -4] exercise
FCPA UPA close out ops
ARED – cylinder flywheel cleaning
SOZh maintenance
EMU SS ion-lithium battery (LIB) recharge (term)
USND2- equipment deactivation and close out ops
EVA procedures review
Audio level measurement with EARQ s/w
DB – experiment run, part 2
Medical operations practicing (subject)
WRS – tank refill from EDV for processing
MELFI 1 Cold Box inspection
ARED cables lubrication
CASCADE. Manual mixing in bio reactor
Exercise data d/l
ANISO – samples stow
Completed Task List Items

Ground Activities
All activities have been completed unless otherwise noted.

UPA activation following UPA troubleshooting
SSRMS maneuver to start position for EVA
Battery Capacity Test on 2A3 and 4A3
Three-Day Look Ahead:

Tuesday, 02/24: EVA Preps, EVA Tool Config, PWD Filter R&R, EVA DOUG and Procedure Reviews and Conference
Wednesday, 02/25: EVA #30 IDA Cable / Lee Lube
Thursday, 02/26: EVA Debrief, EVA Tool Configuration, JEMAL depressurization
QUICK ISS Status – Environmental Control Group:

Component – Status
Elektron – On
Vozdukh – On
[???] 1 – SM Air Conditioner System (“SKV1”) – On
[???] 2 – SM Air Conditioner System (“SKV2”) – Off
Carbon Dioxide Removal Assembly (CDRA) Lab – Standby
Carbon Dioxide Removal Assembly (CDRA) Node 3 – Operate
Major Constituent Analyzer (MCA) Lab – Idle
Major Constituent Analyzer (MCA) Node 3 – Operate
Oxygen Generation Assembly (OGA) – Standby
Urine Processing Assembly (UPA) – Standby
Trace Contaminant Control System (TCCS) Lab – Off
Trace Contaminant Control System (TCCS) Node 3 – Full up

SpaceRef co-founder, entrepreneur, writer, podcaster, nature lover and deep thinker.